# Linear-Size $\mathrm{QAC}^{0}$ Channels: Learning, Testing and Hardness

### v1

@@ Abstract
Shallow quantum circuits have attracted increasing attention in recent years, due to the fact that current noisy quantum hardware can only perform faithful quantum computation for a short amount of time. The constant-depth quantum circuits $\mathbf{QAC}^{0}$, a quantum counterpart of $\mathbf{AC}^{0}$ circuits, are the polynomial-size and constant-depth quantum circuits composed of only single-qubit unitaries and polynomial-size generalized Toffoli gates. The computational power of $\mathbf{QAC}^{0}$ has been extensively investigated in recent years [FFG+06, PFGT20, Ros21, NPVY24, ADOY25].
@@
  • We show that depth-$d$ $\mathbf{QAC}^{0}$ circuits working on $n$ input qubits and $a$ ancilla qubits have approximate degree at most $\tilde{O}\left((n+a)^{1-2^{-d}}\right)$, improving the $\tilde{O}\left((n+a)^{1-3^{-d}}\right)$ degree upper bound of [ADOY25]. Consequently, this directly implies that to compute the parity function, $\mathbf{QAC}^{0}$ circuits need at least $\tilde{O}\left(n^{1+2^{-d}}\right)$ circuit size. We obtain this bound by improving the techniques in [ADOY25] by using the techniques of unitary dilation and operator dilation.
@@ 1 Introduction
Current quantum hardware suffers from decoherence and is only able to perform faithful quantum computation for a short amount of time. This motivates the study of shallow quantum circuits, which have received increasing attention in recent years [BGL24, HLB+24, VH25, ADEGP24]. A simple yet important class of shallow quantum circuits are the $\mathbf{QNC}^{0}$ circuits, which are polynomial size, constant depth quantum circuits with single-qubit and two-qubit quantum gates. Despite these constrains, $\mathbf{QNC}^{0}$ circuits already demonstrate computational advantages over classical algorithms in specific tasks [BGK18, WKST19, BGKT20, WP23], such as sampling from classically intractable distributions [GWD17, HHB+20, CC22]. However, due to the light-cone constraint, the ability of $\mathbf{QNC}^{0}$ circuits to generate long-range entanglement or to compute Boolean functions is significantly limited.
Green, Homer, Moore, and Pollett [Moo99, GHMP01] introduced $\mathbf{QAC}^{0}$ circuits as a quantum counterpart of classical $\mathbf{AC}^{0}$ circuits, which are constant-depth polynomial-size quantum circuits consisting of single-qubit unitary gates together with arbitrarily size generalized Toffoli gates. Classical $\mathbf{AC}^{0}$ is a central object in circuit complexity, whose computational power is well understood and which delineates the frontier of current lower bound techniques. Thus, it is intriguing to investigate the computational power of $\mathbf{QAC}^{0}$. Recent progress on quantum hardware has also enabled the implementation of long-range multi-qubit operations, including the generalized Toffoli gate [RGG+20, GF21, NZB+25] and the quantum fan-out gate [GKH+20], underscoring the relevance of $\mathbf{QAC}^{0}$ as a model for near-term quantum computation.
Using the Pauli analysis framework, Nadimpalli, Parham, Vasconcelos, and Yuen [NPVY24] characterized the structure of $\mathbf{QAC}^{0}$ circuits of depth $d$ and ancilla size $n^{O(1/d)}$. They showed that $\mathbf{QAC}^{0}$ circuits exhibit a low degree Pauli concentration similar to $\mathbf{AC}^{0}$ circuits when the ancilla is of size $n^{O(1/d)}$. Later, Anshu, Dong, Ou, and Yao [ADOY25] demonstrated that $\mathbf{QAC}^{0}$ circuits with barely linear-size $n^{1+3^{-d}}$ have a small approximation degree. Thus, linear-size $\mathbf{QAC}^{0}$ circuits stand at the frontier of the lower bounds of quantum circuits. This motivates us to investigate the family of quantum circuits $\mathbf{QLC}^{0}$  as a quantum analog of $\mathbf{LC}^{0}$ [KPLT06], which are linear size $\mathbf{QAC}^{0}$ circuits.
@@
A central challenge in studying $\mathbf{QAC}^{0}$ circuits is to understand the computational power of $\mathbf{QAC}^{0}$. In particular, given that $\mathbf{AC}^{0}$ circuits can not compute the parity function [Hås86], a core problem is whether $\mathbf{QAC}^{0}$ circuits are stronger than $\mathbf{AC}^{0}$ circuits and can compute the parity function. This problem has been extensively studied in the past [FFG+06, PFGT20, Ros21, NPVY24, ADOY25], where they gave different ancilla lower bounds for $\mathbf{QAC}^{0}$ circuits to compute the parity function. Notably, given that the parity function has linear Fourier degree, the low-degree approximation results of [NPVY24] and [ADOY25] show that $\mathbf{QAC}^{0}$ circuits can not compute the parity function, if there is only $O\left(n^{1/d}\right)$ and $O\left(n^{1+3^{-d}}\right)$ ancilla, respectively. Given that $\mathbf{QAC}^{0}$ circuits with exponential ancilla can compute the parity function [Ros21], a still open question is what is the minimal number of ancilla qubits required to compute the parity function.
An observation of [ADOY25] shows that any ancilla lower bound of the form $n^{1+\exp\left(-o(d)\right)}$ can be boosted to $n^{c}$ for any constant $c>1$. Thus if we can slightly improve the ancilla lower bound of [ADOY25], then we can prove that any $\mathbf{QAC}^{0}$ circuit can not compute the parity function. So one possible path to resolve this open question is to improve the exponent in the $O\left(n^{1+3^{-d}}\right)$ lower bound.
@@
The task of learning an unknown quantum process, a.k.a. Quantum process tomography, is a fundamental task in quantum physics and quantum computation [CN97, PCZ97, DLP01, OPG+04, Sco08, MRL08, HKOT23]. This problem has been extensively studied since the early days of quantum computation and has found broad applications in diverse areas, including quantum cryptography [ADDK21, SHH25], quantum metrology [GLM06, GLM11], and error mitigation [CBB+23].
@@
Haah, Kothari, O’Donnell, and Tang [HKOT23] have shown that learning an arbitrary quantum channel requires exponential resources. However, in practice we are not always faced with learning completely arbitrary quantum channels. Recently a number of learning algorithms for shallow quantum circuits have emerged in the literature [HLB+24, VH25, ADEGP24, BEG24], alongside many learning algorithms for quantum channels with other specific underlying structures, such as junta channels [CNY23, BY23], and Pauli channels [HFW20, FW20, HYF21, CZSJ22, KTCT23, WLKD24]. In the standard model of PAC learning, it assumes that the target channels exactly satisfy the assumed structure. In real-world quantum systems, this assumption is often unrealistic due to unavoidable noise and imperfect gate operations. This motivates the study of agnostic learning, where the learning algorithm does not assume the target lies exactly within the hypothesis class.
@@
Agnostic PAC learning, introduced by Kearns, Schapire, and Sellie [KSS92], is a more general and robust learning model. It seeks to find a hypothesis within a given class that best approximates an arbitrary target function. Agnostic learning has received significant attention since it was introduced. Despite its flexibility, designing agnostic learning algorithms with nontrivial performance guarantees remains a significant challenge, even when allowing sub-exponential running time [BT22].
@@
The quantum version of agnostic learning for functions was first investigated by Arunachalam and de Wolf [ADW18], who established tight lower bounds on the sample complexity of agnostic learning in terms of the VC dimension. More recently, agnostic learning has been extended to quantum states, including product states [BBK+25], and stabilizer states [CGYZ25, GIKL25]. Wadhawa, Lewis, Kashefi, and Doosti [WLKD24] further introduced agnostic process tomography, proposing algorithms for learning several classes of quantum channels, including bounded-gate circuits, Clifford circuits, Pauli strings, $k$-juntas, and low-degree circuits.
@@
Despite this growing body of work, relatively little is known about agnostic learning for shallow quantum circuits. To the best of our knowledge, the only existing result prior to ours in this direction is due to Wadhawan, Lewis, Kashefi, and Doosti [WLKD24] who initiated the study of agnostic learning in this setting. We summarize all currently known results on learning shallow quantum circuits with many-qubit gates, including both upper and lower bounds on their agnostic and non-agnostic learnability in Table 1.
@@
Property testing concerns the task of determining whether a given object satisfies a specified property or is far from satisfying it. It is a central area of research in theoretical computer science. Compared with learning, property testing often admits substantially smaller query complexity. In quantum computing, designing efficient testers for various properties has been studied extensively, [MKasB05, HM13], including for stabilizer states [AD25], junta states [BEG24], quantum junta channels [BY23, CLL24, BLY+25] etc. For an overview, see the survey of Montanaro and de Wolf [MdW13].
The standard notion of property testing requires an algorithm to distinguish between objects that exactly satisfy the property and those that are far from it. However, this notion lacks robustness to noise. To address this, Parnas, Ron, and Rubinfeld [PRR06] introduced tolerant property testing, where the goal is to distinguish objects that approximately satisfy the property from those that are far from it. Tolerant testing is particularly significant in the quantum setting, where noise is inherent and unavoidable. Nonetheless, designing efficient tolerant testers is considerably more challenging and, in some cases, provably impossible [NP24].
@@ 1.1 Our Results
First, similar to [ADOY25], we bound the approximate degree of the output operator of $\mathbf{QLC}^{0}$ channels under low-degree inputs. Specifically, in Section 3, we prove the following theorem:
@@
Our proof leverages the unitary dilation technique to ensure that the approximation of large quantum gates is unitary. By embedding the non-unitary, approximated operator into a larger unitary matrix, the technique ensures that the operator can still be managed and analyzed as part of a valid unitary evolution. This allows us to simultaneously apply the low-degree approximation and light-cone argument to a large generalized Toffoli gate. This critical property leads to a tighter bound on the approximate degree than the one established in [ADOY25].
Compared to the $\tilde{\mathcal{O}}\left(n^{1-3^{d}}\ell^{-3^{d}}\right)$ degree upper bound given by [ADOY25], our degree upper bound of $M$ is slightly better. This result directly implies that the Boolean functions computed by $\mathbf{QLC}^{0}$ circuits also have an approximate degree upper bound of $\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)$, which matches exactly with the approximate degree upper bound for $\mathbf{LC}^{0}$ circuits, thereby demonstrating the existence of a subexponential time agnostic learning algorithm for $\mathbf{QLC}^{0}$ Boolean functions [BKT19].
@@
Our learning algorithm builds upon the framework of Pauli analysis, which involves learning the Pauli coefficients of the Choi state of the quantum channel. Pauli analysis was introduced by Montanaro and Osborne [MO10], and has since become a widely used tool in the design of quantum learning algorithms [CNY23, BY23, NPVY24, ADEGP24, ADEGP25]. This approach combines with classical shadows [HKP20] and Fourier analysis [O’D14], making it particularly effective in scenarios where the Pauli expansion is sparse—such as juntas and low-degree functions.
@@
These hardness results also answer a question raised in [WLKD24], demonstrating that agnostic learning of $\mathbf{QLC}^{0}$ channels is impossible under certain stronger norms.
@@ 1.2 Proof Overview
We combine the technique of unitary dilation and operator dilation with the low-degree approximation results from [ADOY25]. The elementary gates in a $\mathbf{QAC}^{0}$ circuit are the multi-qubit CZ-gates, along with all single-qubit unitaries. The single-qubit unitaries are already of degree $1$, so we do not need further action on them. For large CZ-gate acting on $n$ qubits, it can be approximated by an operator with Pauli degree at most $\sqrt{n}$, up to logarithmic factors.
Here comes the dilation technique: The unitary dilation technique allows us to embed an arbitrary operator with bounded spectral norm into a unitary matrix. This is especially useful when handling low-degree approximations of large quantum gates, which lose the property of being a unitary matrix because of the low-degree approximation. We embed such low-degree approximations into larger unitary operators, and use a layer-by-layer argument to prove an approximate-degree upper bound for $\mathbf{QLC}^{0}$ circuits, using similar ideas as in [ADOY25].
@@
So to agnostically learn an arbitrary operator $P$, we only need to learn the low-degree part $P^{\leq d}$, which consists of at most $(3n)^{d}$ Pauli coefficients. This can be achieved efficiently using classical shadow tomography from Huang, Kueng, and Preskill [HKP20, Lemma 17].
@@
In the first step, we use a reduction from channel hardness to unitary hardness. The main ingredient of the reduction is similar to that in the algorithm [VH25, Algorithm 1] provided by Vasconcelos and Huang. For a $\mathbf{QAC}^{0}$ unitary $V$ with $a$ ancilla, suppose $V_{i}$ is the circuit related to the $i$-th qubit in the sense that $\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)=\mathrm{Tr}_{-i}\>\!\!\left(V_{i}\rho V_{i}^{\dagger}\right)$. Broadly speaking, the beauty of this algorithm lies in the fact that, using the local information of $V_{i}$, we can sew together the global $V\otimes V^{\dagger}$. Conversely, as a reduction, if learning the global unitary $V\otimes V^{\dagger}$ is difficult, then learning the local $V_{i}$ should also be difficult.
To prove the hardness of learning $\mathbf{QAC}^{0}$ $n\to 1$ channels, we now only need the hardness of learning $V\otimes V^{\dagger}$. This is done with a slightly tailored result from [VH25, Proposition 7].

### v2

@@ Abstract
Shallow quantum circuits have attracted increasing attention in recent years, due to the fact that current noisy quantum hardware can only perform faithful quantum computation for a short amount of time. The constant-depth quantum circuits $\mathbf{QAC}^{0}$, a quantum counterpart of $\mathbf{AC}^{0}$ circuits, are the polynomial-size and constant-depth quantum circuits composed of only single-qubit unitaries and polynomial-size generalized Toffoli gates. The computational power of $\mathbf{QAC}^{0}$ has been extensively investigated in recent years [10.5555/2011679.2011682, DBLP:journals/corr/abs-2005-12169, rosenthal:LIPIcs.ITCS.2021.32, NPVY24, ADOY24].
@@
  • We show that depth-$d$ $\mathbf{QAC}^{0}$ circuits working on $n$ input qubits and $a$ ancilla qubits have approximate degree at most $\tilde{O}\left((n+a)^{1-2^{-d}}\right)$, improving the $\tilde{O}\left((n+a)^{1-3^{-d}}\right)$ degree upper bound of [ADOY24]. Consequently, this directly implies that to compute the parity function, $\mathbf{QAC}^{0}$ circuits need at least $\tilde{O}\left(n^{1+2^{-d}}\right)$ circuit size. We obtain this bound by improving the techniques in [ADOY24] by using the techniques of unitary dilation and operator dilation.
@@ 1 Introduction
Current quantum hardware suffers from decoherence and is only able to perform faithful quantum computation for a short amount of time. This motivates the study of shallow quantum circuits, which have received increasing attention in recent years [bravyi2024classical, HLB+24, vasconcelos2024learning, arunachalam_et_al:LIPIcs.ICALP.2024.13]. A simple yet important class of shallow quantum circuits are the $\mathbf{QNC}^{0}$ circuits, which are polynomial size, constant depth quantum circuits with single-qubit and two-qubit quantum gates. Despite these constrains, $\mathbf{QNC}^{0}$ circuits already demonstrate computational advantages over classical algorithms in specific tasks [doi:10.1126/science.aar3106, 10.1145/3313276.3316404, bravyi2020quantum, watts2023unconditional], such as sampling from classically intractable distributions [gao2017quantum, haferkamp2020closing, coble2022quasi]. However, due to the light-cone constraint, the ability of $\mathbf{QNC}^{0}$ circuits to generate long-range entanglement or to compute Boolean functions is significantly limited.
Green, Homer, Moore, and Pollett [moore1999quantum, moor] introduced $\mathbf{QAC}^{0}$ circuits as a quantum counterpart of classical $\mathbf{AC}^{0}$ circuits, which are constant-depth polynomial-size quantum circuits consisting of single-qubit unitary gates together with arbitrarily size generalized Toffoli gates. Classical $\mathbf{AC}^{0}$ is a central object in circuit complexity, whose computational power is well understood and which delineates the frontier of current lower bound techniques. Thus, it is intriguing to investigate the computational power of $\mathbf{QAC}^{0}$. Recent progress on quantum hardware has also enabled the implementation of long-range multi-qubit operations, including the generalized Toffoli gate [rasmussen2020single, goel2021native, nikolaeva2025scalable] and the quantum fan-out gate [Gokhale2020QuantumFC], underscoring the relevance of $\mathbf{QAC}^{0}$ as a model for near-term quantum computation.
Using the Pauli analysis framework, Nadimpalli, Parham, Vasconcelos, and Yuen [NPVY24] characterized the structure of $\mathbf{QAC}^{0}$ circuits of depth $d$ and ancilla size $n^{O(1/d)}$. They showed that $\mathbf{QAC}^{0}$ circuits exhibit a low degree Pauli concentration similar to $\mathbf{AC}^{0}$ circuits when the ancilla is of size $n^{O(1/d)}$. Later, Anshu, Dong, Ou, and Yao [ADOY24] demonstrated that $\mathbf{QAC}^{0}$ circuits with barely linear-size $n^{1+3^{-d}}$ have a small approximation degree. Thus, linear-size $\mathbf{QAC}^{0}$ circuits stand at the frontier of the lower bounds of quantum circuits. This motivates us to investigate the family of quantum circuits $\mathbf{QLC}^{0}$  as a quantum analog of $\mathbf{LC}^{0}$ [1663737], which are linear size $\mathbf{QAC}^{0}$ circuits.
@@
A central challenge in studying $\mathbf{QAC}^{0}$ circuits is to understand the computational power of $\mathbf{QAC}^{0}$. In particular, given that $\mathbf{AC}^{0}$ circuits can not compute the parity function [10.1145/12130.12132], a core problem is whether $\mathbf{QAC}^{0}$ circuits are stronger than $\mathbf{AC}^{0}$ circuits and can compute the parity function. This problem has been extensively studied in the past [10.5555/2011679.2011682, DBLP:journals/corr/abs-2005-12169, rosenthal:LIPIcs.ITCS.2021.32, NPVY24, ADOY24], where they gave different ancilla lower bounds for $\mathbf{QAC}^{0}$ circuits to compute the parity function. Notably, given that the parity function has linear Fourier degree, the low-degree approximation results of [NPVY24] and [ADOY24] show that $\mathbf{QAC}^{0}$ circuits can not compute the parity function, if there is only $O\left(n^{1/d}\right)$ and $O\left(n^{1+3^{-d}}\right)$ ancilla, respectively. Given that $\mathbf{QAC}^{0}$ circuits with exponential ancilla can compute the parity function [rosenthal:LIPIcs.ITCS.2021.32], a still open question is what is the minimal number of ancilla qubits required to compute the parity function.
An observation of [ADOY24] shows that any ancilla lower bound of the form $n^{1+\exp\left(-o(d)\right)}$ can be boosted to $n^{c}$ for any constant $c>1$. Thus if we can slightly improve the ancilla lower bound of [ADOY24], then we can prove that any $\mathbf{QAC}^{0}$ circuit can not compute the parity function. So one possible path to resolve this open question is to improve the exponent in the $O\left(n^{1+3^{-d}}\right)$ lower bound.
@@
The task of learning an unknown quantum process, a.k.a. Quantum process tomography, is a fundamental task in quantum physics and quantum computation [Chuang_1997, Poyatos_1997, PhysRevLett.86.4195, O_Brien_2004, Scott_2008, PhysRevA.77.032322, Haah_2023]. This problem has been extensively studied since the early days of quantum computation and has found broad applications in diverse areas, including quantum cryptography [Arapinis2021quantumphysical, doi:10.1126/science.adv8590], quantum metrology [PhysRevLett.96.010401, Giovannetti2011], and error mitigation [RevModPhys.95.045005].
@@
Haah, Kothari, O’Donnell, and Tang [Haah_2023] have shown that learning an arbitrary quantum channel requires exponential resources. However, in practice we are not always faced with learning completely arbitrary quantum channels. Recently a number of learning algorithms for shallow quantum circuits have emerged in the literature [HLB+24, vasconcelos2024learning, arunachalam_et_al:LIPIcs.ICALP.2024.13, bao2024learning], alongside many learning algorithms for quantum channels with other specific underlying structures, such as junta channels [doi:10.1137/1.9781611977554.ch43, pmlr-v195-bao23b], and Pauli channels [Harper_2020, Flammia_2020, Harper_2021, Chen_2022, Kunjummen_2023, wadhwa2024agnostic]. In the standard model of PAC learning, it assumes that the target channels exactly satisfy the assumed structure. In real-world quantum systems, this assumption is often unrealistic due to unavoidable noise and imperfect gate operations. This motivates the study of agnostic learning, where the learning algorithm does not assume the target lies exactly within the hypothesis class.
@@
Agnostic PAC learning, introduced by Kearns, Schapire, and Sellie [10.1145/130385.130424], is a more general and robust learning model. It seeks to find a hypothesis within a given class that best approximates an arbitrary target function. Agnostic learning has received significant attention since it was introduced. Despite its flexibility, designing agnostic learning algorithms with nontrivial performance guarantees remains a significant challenge, even when allowing sub-exponential running time [10.1145/3444815.3444825].
@@
The quantum version of agnostic learning for functions was first investigated by Arunachalam and de Wolf [10.5555/3291125.3309633], who established tight lower bounds on the sample complexity of agnostic learning in terms of the VC dimension. More recently, agnostic learning has been extended to quantum states, including product states [10.1145/3717823.3718207], and stabilizer states [chen2024stabilizerbootstrappingrecipeefficient, grewal2025agnostictomographystabilizerproduct]. Wadhwa, Lewis, Kashefi, and Doosti [wadhwa2024agnostic] further introduced agnostic process tomography, proposing algorithms for learning several classes of quantum channels, including bounded-gate circuits, Clifford circuits, Pauli strings, $k$-juntas, and low-degree circuits.
@@
Despite this growing body of work, relatively little is known about agnostic learning for shallow quantum circuits. To the best of our knowledge, the only existing result prior to ours in this direction is due to Wadhawan, Lewis, Kashefi, and Doosti [wadhwa2024agnostic] who initiated the study of agnostic learning in this setting. We summarize all currently known results on learning shallow quantum circuits with many-qubit gates, including both upper and lower bounds on their agnostic and non-agnostic learnability in Table 1.
@@
Property testing concerns the task of determining whether a given object satisfies a specified property or is far from satisfying it. It is a central area of research in theoretical computer science. Compared with learning, property testing often admits substantially smaller query complexity. In quantum computing, designing efficient testers for various properties has been studied extensively, [PhysRevLett.95.260502, harrow2013testing], including for stabilizer states [10.1145/3717823.3718277], junta states [bao2024learning], quantum junta channels [pmlr-v195-bao23b, chen2024tolerant, bao2025efficient] etc. For an overview, see the survey of Montanaro and de Wolf [montanaro2013survey].
The standard notion of property testing requires an algorithm to distinguish between objects that exactly satisfy the property and those that are far from it. However, this notion lacks robustness to noise. To address this, Parnas, Ron, and Rubinfeld [PARNAS20061012] introduced tolerant property testing, where the goal is to distinguish objects that approximately satisfy the property from those that are far from it. Tolerant testing is particularly significant in the quantum setting, where noise is inherent and unavoidable. Nonetheless, designing efficient tolerant testers is considerably more challenging and, in some cases, provably impossible [10.1145/3618260.3649687].
@@ 1.1 Our Results
First, similar to [ADOY24], we bound the approximate degree of the output operator of $\mathbf{QLC}^{0}$ channels under low-degree inputs. Specifically, in Section 3, we prove the following theorem:
@@
Our proof leverages the unitary dilation technique to ensure that the approximation of large quantum gates is unitary. By embedding the non-unitary, approximated operator into a larger unitary matrix, the technique ensures that the operator can still be managed and analyzed as part of a valid unitary evolution. This allows us to simultaneously apply the low-degree approximation and light-cone argument to a large generalized Toffoli gate. This critical property leads to a tighter bound on the approximate degree than the one established in [ADOY24].
Compared to the $\tilde{\mathcal{O}}\left(n^{1-3^{d}}\ell^{-3^{d}}\right)$ degree upper bound given by [ADOY24], our degree upper bound of $M$ is slightly better. This result directly implies that the Boolean functions computed by $\mathbf{QLC}^{0}$ circuits also have an approximate degree upper bound of $\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)$, which matches exactly with the approximate degree upper bound for $\mathbf{LC}^{0}$ circuits, thereby demonstrating the existence of a subexponential time agnostic learning algorithm for $\mathbf{QLC}^{0}$ Boolean functions [doi:10.1137/1.9781611975482.42].
@@
Our learning algorithm builds upon the framework of Pauli analysis, which involves learning the Pauli coefficients of the Choi state of the quantum channel. Pauli analysis was introduced by Montanaro and Osborne [MO10], and has since become a widely used tool in the design of quantum learning algorithms [doi:10.1137/1.9781611977554.ch43, pmlr-v195-bao23b, NPVY24, arunachalam_et_al:LIPIcs.ICALP.2024.13, Arunachalam2025]. This approach combines with classical shadows [Huang_2020] and Fourier analysis [ODonnell2014], making it particularly effective in scenarios where the Pauli expansion is sparse—such as juntas and low-degree functions.
@@
These hardness results also answer a question raised in [wadhwa2024agnostic], demonstrating that agnostic learning of $\mathbf{QLC}^{0}$ channels is impossible under certain stronger norms.
@@ 1.2 Proof Overview
We combine the technique of unitary dilation and operator dilation with the low-degree approximation results from [ADOY24]. The elementary gates in a $\mathbf{QAC}^{0}$ circuit are the multi-qubit CZ-gates, along with all single-qubit unitaries. The single-qubit unitaries are already of degree $1$, so we do not need further action on them. For large CZ-gate acting on $n$ qubits, it can be approximated by an operator with Pauli degree at most $\sqrt{n}$, up to logarithmic factors.
Here comes the dilation technique: The unitary dilation technique allows us to embed an arbitrary operator with bounded spectral norm into a unitary matrix. This is especially useful when handling low-degree approximations of large quantum gates, which lose the property of being a unitary matrix because of the low-degree approximation. We embed such low-degree approximations into larger unitary operators, and use a layer-by-layer argument to prove an approximate-degree upper bound for $\mathbf{QLC}^{0}$ circuits, using similar ideas as in [ADOY24].
@@
So to agnostically learn an arbitrary operator $P$, we only need to learn the low-degree part $P^{\leq d}$, which consists of at most $(3n)^{d}$ Pauli coefficients. This can be achieved efficiently using classical shadow tomography from Huang, Kueng, and Preskill [Huang_2020, Lemma 17].
@@
In the first step, we use a reduction from channel hardness to unitary hardness. The main ingredient of the reduction is similar to that in the algorithm [vasconcelos2024learning, Algorithm 1] provided by Vasconcelos and Huang. For a $\mathbf{QAC}^{0}$ unitary $V$ with $a$ ancilla, suppose $V_{i}$ is the circuit related to the $i$-th qubit in the sense that $\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)=\mathrm{Tr}_{-i}\>\!\!\left(V_{i}\rho V_{i}^{\dagger}\right)$. Broadly speaking, the beauty of this algorithm lies in the fact that, using the local information of $V_{i}$, we can sew together the global $V\otimes V^{\dagger}$. Conversely, as a reduction, if learning the global unitary $V\otimes V^{\dagger}$ is difficult, then learning the local $V_{i}$ should also be difficult.
To prove the hardness of learning $\mathbf{QAC}^{0}$ $n\to 1$ channels, we now only need the hardness of learning $V\otimes V^{\dagger}$. This is done with a slightly tailored result from [vasconcelos2024learning, Proposition 7].